Two MRT trains leave Guadalupe stations at the same time, traveling in opposite direction. The speed of one train is 5mph faster than the other train. At the end of 30 minutes, they are 6 miles apart. Find the rates of speed for each train.
tnx for answering in advance!!
: Two MRT trains leave Guadalupe stations at the same time, traveling in opposite direction. The speed of one train is 5mph faster than the other train. At the end of 30 minutes, they are 6 miles apart. Find the rates of speed for each train.
tnx for answering in advance!!

Answer by galileo(7)

Let x = speed of 1st train
x+5 = speed of 2nd train
After 30 minutes the distance between the two trains is equal to the sum of their distances traveled.
x(0.5) + (x+5)(0.5) = 6
(0.5)(x+x+5) = 6 ( By Factoring )
(0.5)(2x+5) = 6
2x+5 = 6/0.5 ( Eliminating 0.5 )
2x+5 = 12
2x = 12-5 ( Transposition )
2x = 7
x = 7/2 ( Divide both sides by 2 )
x = 3.5 mph (speed of 1st train)
x+5 = 8.5 mph (speed of 2nd train)